OpenStudy (anonymous):

A test of a new car results in 540 miles on 10 gallons of gas. How far could you drive on 45 gallons of gas? What is the car's gas mileage?

OpenStudy (studygurl14):

540 / 10 = miles per gallon multiply that answer ^ by 45 to get how far you could drive on 45 gallons

OpenStudy (anonymous):

540/2 = 270 540*4 = 2160 2160 + 270 = 2430